OpenStudy (anonymous):

What is the value of the expression? 4^3 • 4^−5

OpenStudy (shamim):

x^3*x^7=?

OpenStudy (anonymous):

4^-15?

OpenStudy (shamim):

No

OpenStudy (anonymous):

A. 4^−15 B. 4^−2 C. 4^2 D. 4^8

OpenStudy (anonymous):

d

OpenStudy (shamim):

4^(3-5)=?

563blackghost (563blackghost):

you add the exponents

563blackghost (563blackghost):

and the base stays the same so whats the answer?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

4^-2

563blackghost (563blackghost):

So (3 + -5) is -2 so the answer is B. xD

OpenStudy (shamim):

3-5=?

OpenStudy (anonymous):

can you help with another